# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v3

@@ Abstract
Differentially Private Stochastic Gradient Descent (DPSGD) is widely used to protect sensitive data during the training of machine learning models, but its privacy guarantee often comes at a large cost of model performance due to the lack of tight theoretical bounds quantifying privacy loss. While recent efforts have achieved more accurate privacy guarantees, they still impose some assumptions prohibited from practical applications, such as convexity and complex parameter requirements, and rarely investigate in-depth the impact of privacy mechanisms on the model’s utility. In this paper, we provide a rigorous privacy characterization for DPSGD with general $L$-smooth and non-convex loss functions, revealing converged privacy loss with iteration in bounded-domain cases. Specifically, we track the privacy loss over multiple iterations, leveraging the noisy smooth-reduction property, and further establish comprehensive convergence analysis in different scenarios. In particular, we show that for DPSGD with a bounded domain, (i) the privacy loss can still converge without the convexity assumption, (ii) a smaller bounded diameter can improve both privacy and utility simultaneously under certain conditions, and (iii) the attainable big-$\mathcal{O}$ order of the privacy utility trade-off for DPSGD with gradient clipping (DPSGD-GC) and for DPSGD-GC with bounded domain (DPSGD-DC) and $\mu$-strongly convex population risk function, respectively. Experiments via membership inference attack (MIA) in a practical setting validate insights gained from the theoretical results.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DPSGD) (abadi2016deep) has emerged as the leading defense mechanism to protect personal sensitive data in training of machine learning models. However, achieving good performance with DPSGD often comes with a significant privacy cost. A fundamental question, therefore, is how to precisely quantify the privacy loss associated with DPSGD.
Previous methods for quantifying privacy loss include strong composition (dwork2010boosting; bassily2014private; kairouz2015composition), moments accountant (abadi2016deep), Rényi Differential Privacy (RDP) (mironov2017renyi; mironov2019r), and Gaussian Differential Privacy (GDP) (dong2022gaussian), along with several numerical composition methods (koskela2020computing; gopi2021numerical). These methods primarily rely on composition theorems, assuming that all intermediate models are revealed during the training procedure, which leads to an overestimation of privacy loss. While numerical composition methods aim to tightly characterize the privacy loss, they still operate under this same assumption.
To address this overestimation, recent works have focused solely on the privacy guarantees of the final output. For instance, the privacy amplification by iteration (feldman2018privacy) demonstrated that withholding intermediate results significantly enhances privacy guarantees for smooth and convex objectives. Building upon this, chourasia2021differential suggest that the privacy loss of DPGD, the full batch version of DPSGD, may converge exponentially fast for smooth and strongly convex objectives. Furthermore, results by ye2022differentially as well as ryffel2022differential extended this analysis to assess the privacy loss of DPSGD, although both studies rely on the assumption of strong convexity.
More recently, the work by altschuler2022privacy and its extension (altschuler2024privacy) established a constant upper bound on privacy loss after a burn-in period for Lipschitz continuous and smooth convex losses over a bounded domain. However, this analytical result is limited by its reliance on the convexity assumption and strict restrictions on the Rényi parameter $\alpha$, which hinders its broader applicability. In order to relax several strong assumptions, kong2024privacy provided an analysis of weakly-convex smooth losses in the case where data is traversed cyclically. Later, chien2024convergent suggest precisely tracking the privacy leakage incurred before reaching the constant upper bound by solving an optimization problem. However, this result is formulated as a complex optimization problem rather than a closed-form expression, making it hard to operationalize. Notably, most recent methods necessitate double clipping of both gradients and parameters due to the additional bounded domain assumption. These methods, however, do not provide a thorough utility analysis or experimental results, leaving their practical performance and trade-offs underexplored.
@@
feldman2018privacy convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ unbounded $\mathcal{O}\left(\frac{\alpha M^{2}}{b^{2}\sigma_{\text{DP}}^{2}}T\right)$ $\surd$
altschuler2022privacy convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ bounded $\mathcal{O}\left(\frac{\alpha M^{2}}{n^{2}\sigma_{\text{DP}}^{2}}\min\left\{T,\frac{Dn}{\eta M}\right\}\right)$ $\times$
@@
kong2024privacy $m$-weakly convex, $L$-smooth, $\eta\leq\frac{1}{2(m+L)}$ bounded $\mathcal{O}\left(\frac{\alpha}{\eta^{2}\sigma_{\text{DP}}^{2}}(D\sqrt{1+2\eta m[1+\frac{m}{2(L+m)}]}+\frac{\eta C}{b})^{2}\right)$ $\times$
chien2024convergent $M$-Lipschitz unbounded w/o analytical form $\times$
@@ 1.1 Contributions
To clearly illustrate the effectiveness of our proposed privacy bound, we here provide detailed comparisons with several prominent existing approaches, including: a) feldman2018privacy, (b) the combined analysis by mironov2017renyi and mironov2019r, c) altschuler2022privacy, and d) kong2024privacy. The detailed setting can be found in Appendix D.1. As shown in Figure 1, our analysis demonstrates strict improvement on all existing privacy bounds, except for altschuler2022privacy. The reason for which their bound appeared tighter is their additional assumptions—including convexity and more restrictive conditions (as summarized in Table 1)—while our analysis relies only on the smoothness of the loss function and far weaker assumptions.
@@ 1.2 Other Related Works
In 2014, bassily2014private analyzed the optimal utility guarantees of DPSGD under the assumption of Lipschitz continuity, considering both convex and strongly convex cases. Then, based on the additional assumption of the gradient distribution, chen2020understanding studied the convergence of DPSGD with gradient clipping (DPSGD-GC) and derived a utility bound for the non-convex setting. The work by song2021evading explored the convergence of DPSGD-GC for generalized linear models, noting that, in the worst case, the utility can remain constant relative to the original objective. Later, fang2023improved further refined this analysis for smooth and unconstrained problems, providing more precise convergence results. However, many of these studies fix a specific value for the clipping threshold $C$, which may be adjusted due to privacy requirements.
More recently, koloskova2023revisiting characterized the convergence guarantees for DPSGD-GC across various clipping thresholds $C$ in the non-convex setting. While this work provides valuable convergence insights for DPSGD-GC, recent privacy characterizations have introduced the need for double clipping—clipping both gradients and parameters—due to the additional assumption of bounded domains. The convergence analysis involving double clipping has not been thoroughly explored in the existing literature.
@@ 2.1 Rényi Differential Privacy (RDP)
(Differential privacy (dwork2006our)). For $\epsilon\geq 0$, $\delta\in[0,1]$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is $(\epsilon,\delta)$-DP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, and for any subset of outputs $\mathcal{S}\subseteq\mathcal{Y}$, we have
@@
(Rényi divergence (renyi1961measures)). For adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$, and an outcome $s\in\mathcal{Y}$, the Rényi divergence of a finite order $\alpha\neq 1$ between $\mathcal{M}(\mathcal{D})$ and $\mathcal{M}(\mathcal{D}^{\prime})$ is defined as
@@
(Rényi differential privacy (mironov2017renyi)). For $\alpha>1$, $\varepsilon\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ satisfies $(\alpha,\varepsilon)$-RDP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, it holds that
@@
(From $(\alpha,\varepsilon)$-RDP to $(\epsilon,\delta)$-DP (mironov2017renyi)). If $\mathcal{M}$ is an $(\alpha,\varepsilon)$-RDP mechanism, it is also $(\varepsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $0<\delta<1$.
Based on the assumption that intermediate training models are not revealed, the technique privacy amplification by iteration substantially improves the privacy guarantee analysis (feldman2018privacy), which is grounded on the concept of shifted Rényi divergence, as follows.
@@
(Shifted Rényi divergence (feldman2018privacy)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $z\geq 0$ and $\alpha>1$, the $z$-shifted Rényi divergence is defined as
@@
(Shift-reduction (feldman2018privacy)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $a\geq 0$ and $z\geq 0$, we have
@@ 3.1 Privacy Analysis of DPSGD
Note that this result generalizes the contraction-reduction lemma (see Lemma A.1) (feldman2018privacy) and its variants (altschuler2022privacy; altschuler2024privacy) in existing literature, which all rely on the convexity of the loss function to ensure that the update function is contractive. [^3] A function is said to be contractive if it is $1$-Lipschitz. It characterizes the privacy dynamics of shifted Rényi divergence for noisy stochastic updates with general smooth loss functions. Based on this building block, we are ready to present the privacy guarantees for DPSGD-GC and DPSGD-DC, respectively.
@@
Comparing Theorem 3.4 to 3.3, we observe that the privacy loss for DGSGD-DC can still converge to a constant for non-convex smooth loss function when the model-parameter domain is bounded by $D$. However, unlike the convex cases (altschuler2022privacy; altschuler2024privacy), where the upper bound scales linearly with $D$, the non-convex setting yields a bound that scales quadratically with $D$. This suggests that upper bounds on RDP for bounded-domain scenarios are inherently looser without the assumption on convexity, which is consistent with intuition and recent findings (kong2024privacy; chien2024convergent).
@@ 3.2 Utility Analysis of DPSGD
(Utility bound for DPSGD-GC (koloskova2023revisiting)). Assume that the population risk function $l(\cdot)$ has smoothness parameter $L$ and SGD variance bounded from above by $\sigma_{\textup{SGD}}^{2}$ (Assumption 3.6). The DPSGD-GC running for $T$ iterations with step-size $\eta\leq\frac{1}{9L}$ has the minimum expected norm of the population risk’s gradient upper bounded by
@@
Unlike previous works on DPSGD-DC (altschuler2022privacy; altschuler2024privacy; kong2024privacy; chien2024convergent), which mainly focused on RDP analysis, our results explicitly demonstrate how gradient clipping and projection affect the utility of DPSGD-DC. As seen from Theorem 3.4, we establish an eventually constant upper bound on privacy loss dependent on the bounded domain diameter $D$, which yields tighter privacy guarantees with smaller $D$. Meanwhile, Proposition 3.12 demonstrates that when the conditions in Theorem 3.10 are satisfied, a smaller $D$ also leads to a lower upper bound on the utility, thus enhancing the privacy-utility trade-off. [^4] A brief discussion on selecting $D$ is provided in Appendix E.3.
@@ 4 Experiment Evaluation
In this section, we present empirical results estimating the privacy level via MIA. Following kairouz2015composition, we estimate $(\epsilon,\delta)$-DP using the false positive rate (FPR) and false negative rate (FNR) of its attack model on the test data, applying the following formula:
@@ 5 Conclusions
Acknowledgments
The authors would like to thank the anonymous reviewers, SPC, and AC for their insightful comments. The work of H. Xing was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2025A1515010123, and in part by the Guangzhou Municipal Science and Technology Project under Grants 2024A04J4527 and 2023A03J0663.
@@ A.1 Privacy Amplification by Iteration
Lemma A.1 (Contraction-reduction (feldman2018privacy)).
@@
Altschuler et al. (altschuler2022privacy; altschuler2024privacy) provide a mild generalization of this lemma, extending it to be suitable for random contraction maps.
@@ A.3 Lemmas for Privacy Analysis
Definition A.7 ($\infty$-Wasserstein distance (villani2009wasserstein)).
@@
Lemma A.8 (Equivalent definitions of $\infty$-Wasserstein distance (feldman2018privacy)).
@@
Lemma A.9 (Post-processing inequality (van2014renyi)).
@@
Lemma A.10 (Partial convexity in the second argument (van2014renyi)).
@@
Lemma A.11 (Strong composition for Rényi divergence (mironov2017renyi)).
@@
Lemma A.13 (Upper bound on Rényi divergence of the SGM (mironov2019r)).

### v4

@@ Abstract
Differentially Private Stochastic Gradient Descent (DPSGD) is widely used to protect sensitive data during the training of machine learning models, but its privacy guarantee often comes at a large cost of model performance due to the lack of tight theoretical bounds quantifying privacy loss. While recent efforts have achieved more accurate privacy guarantees, they still impose some assumptions prohibited from practical applications, such as convexity and complex parameter requirements, and rarely investigate in-depth the impact of privacy mechanisms on the model’s utility. In this paper, we provide a rigorous privacy characterization for DPSGD with general L-smooth and non-convex loss functions, revealing converged privacy loss with iteration in bounded-domain cases. Specifically, we track the privacy loss over multiple iterations, leveraging the noisy smooth-reduction property, and further establish comprehensive convergence analysis in different scenarios. In particular, we show that for DPSGD with a bounded domain, (i) the privacy loss can still converge without the convexity assumption, (ii) a smaller bounded diameter can improve both privacy and utility simultaneously under certain conditions, and (iii) the attainable big-O order of the privacy utility trade-off for DPSGD with gradient clipping (DPSGD-GC) and for DPSGD-GC with bounded domain (DPSGD-DC) and strongly convex population risk function, respectively. Experiments via membership inference attack (MIA) in a practical setting validate insights gained from the theoretical results.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DPSGD) (Abadi et al. 2016) has emerged as the leading defense mechanism to protect personal sensitive data in training of machine learning models. However, achieving good performance with DPSGD often comes with a significant privacy cost. A fundamental question, therefore, is how to precisely quantify the privacy loss associated with DPSGD.
Previous methods for quantifying privacy loss include strong composition (Dwork et al. 2010; Bassily et al. 2014; Kairouz et al. 2015), moments accountant (Abadi et al. 2016), Rényi Differential Privacy (RDP) (Mironov 2017; Mironov et al. 2019), and Gaussian Differential Privacy (GDP) (Dong et al. 2022), along with several numerical composition methods (Koskela et al. 2020; Gopi et al. 2021). These methods primarily rely on composition theorems, assuming that all intermediate models are revealed during the training procedure, which leads to an overestimation of privacy loss. While numerical composition methods aim to tightly characterize the privacy loss, they still operate under this same assumption.
To address this overestimation, recent works have focused solely on the privacy guarantees of the final output. For instance, the privacy amplification by iteration (Feldman et al. 2018) demonstrated that withholding intermediate results significantly enhances privacy guarantees for smooth and convex objectives. Building upon this, Chourasia et al. (2021) suggest that the privacy loss of DPGD, the full batch version of DPSGD, may converge exponentially fast for smooth and strongly convex objectives. Furthermore, results by Ye and Shokri (2022) as well as Ryffel et al. (2022) extended this analysis to assess the privacy loss of DPSGD, although both studies rely on the assumption of strong convexity.
More recently, the work by Altschuler and Talwar (2022) and its extension (Altschuler et al. 2024) established a constant upper bound on privacy loss after a burn-in period for Lipschitz continuous and smooth convex losses over a bounded domain. However, this analytical result is limited by its reliance on the convexity assumption and strict restrictions on the Rényi parameter $\alpha$, which hinders its broader applicability. In order to relax several strong assumptions, Kong and Ribero (2024) provided an analysis of weakly-convex smooth losses in the case where data is traversed cyclically. Later, Chien and Li (2024) suggest precisely tracking the privacy leakage incurred before reaching the constant upper bound by solving an optimization problem. However, this result is formulated as a complex optimization problem rather than a closed-form expression, making it hard to operationalize. Notably, most recent methods necessitate double clipping of both gradients and parameters due to the additional bounded domain assumption. These methods, however, do not provide a thorough utility analysis or experimental results, leaving their practical performance and trade-offs underexplored.
@@
Feldman et al. (2018) convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ unbounded $\mathcal{O}\left(\frac{\alpha M^{2}}{b^{2}\sigma_{\text{DP}}^{2}}T\right)$ $\surd$
Altschuler and Talwar (2022) convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ bounded $\mathcal{O}\left(\frac{\alpha M^{2}}{n^{2}\sigma_{\text{DP}}^{2}}\min\left\{T,\frac{Dn}{\eta M}\right\}\right)$ $\times$
@@
Kong and Ribero (2024) $m$-weakly convex, $L$-smooth, $\eta\leq\frac{1}{2(m+L)}$ bounded $\mathcal{O}\left(\frac{\alpha}{\eta^{2}\sigma_{\text{DP}}^{2}}(D\sqrt{1+2\eta m[1+\frac{m}{2(L+m)}]}+\frac{\eta C}{b})^{2}\right)$ $\times$
Chien and Li (2024) $M$-Lipschitz unbounded w/o analytical form $\times$
@@ 1.1 Contributions
To clearly illustrate the effectiveness of our proposed privacy bound, we here provide detailed comparisons with several prominent existing approaches, including: a) Feldman et al. (2018), (b) the combined analysis by Mironov (2017) and Mironov et al. (2019), c) Altschuler and Talwar (2022), and d) Kong and Ribero (2024). The detailed setting can be found in Appendix D.1. As shown in Figure 1, our analysis demonstrates strict improvement on all existing privacy bounds, except for Altschuler and Talwar (2022). The reason for which their bound appeared tighter is their additional assumptions—including convexity and more restrictive conditions (as summarized in Table 1)—while our analysis relies only on the smoothness of the loss function and far weaker assumptions.
@@ 1.2 Other Related Works
In 2014, Bassily et al. (2014) analyzed the optimal utility guarantees of DPSGD under the assumption of Lipschitz continuity, considering both convex and strongly convex cases. Then, based on the additional assumption of the gradient distribution, Chen et al. (2020) studied the convergence of DPSGD with gradient clipping (DPSGD-GC) and derived a utility bound for the non-convex setting. The work by Song et al. (2021) explored the convergence of DPSGD-GC for generalized linear models, noting that, in the worst case, the utility can remain constant relative to the original objective. Later, Fang et al. (2023) further refined this analysis for smooth and unconstrained problems, providing more precise convergence results. However, many of these studies fix a specific value for the clipping threshold $C$, which may be adjusted due to privacy requirements.
More recently, Koloskova et al. (2023) characterized the convergence guarantees for DPSGD-GC across various clipping thresholds $C$ in the non-convex setting. While this work provides valuable convergence insights for DPSGD-GC, recent privacy characterizations have introduced the need for double clipping—clipping both gradients and parameters—due to the additional assumption of bounded domains. The convergence analysis involving double clipping has not been thoroughly explored in the existing literature.
@@ 2.1 Rényi Differential Privacy (RDP)
(Differential privacy (Dwork et al. 2006)). For $\epsilon\geq 0$, $\delta\in[0,1]$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is $(\epsilon,\delta)$-DP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, and for any subset of outputs $\mathcal{S}\subseteq\mathcal{Y}$, we have
@@
(Rényi divergence (Rényi 1961)). For adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$, and an outcome $s\in\mathcal{Y}$, the Rényi divergence of a finite order $\alpha\neq 1$ between $\mathcal{M}(\mathcal{D})$ and $\mathcal{M}(\mathcal{D}^{\prime})$ is defined as
@@
(Rényi differential privacy (Mironov 2017)). For $\alpha>1$, $\varepsilon\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ satisfies $(\alpha,\varepsilon)$-RDP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, it holds that
@@
(From $(\alpha,\varepsilon)$-RDP to $(\epsilon,\delta)$-DP (Mironov 2017)). If $\mathcal{M}$ is an $(\alpha,\varepsilon)$-RDP mechanism, it is also $(\varepsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $0<\delta<1$.
Based on the assumption that intermediate training models are not revealed, the technique privacy amplification by iteration substantially improves the privacy guarantee analysis (Feldman et al. 2018), which is grounded on the concept of shifted Rényi divergence, as follows.
@@
(Shifted Rényi divergence (Feldman et al. 2018)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $z\geq 0$ and $\alpha>1$, the $z$-shifted Rényi divergence is defined as
@@
(Shift-reduction (Feldman et al. 2018)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $a\geq 0$ and $z\geq 0$, we have
@@ 3.1 Privacy Analysis of DPSGD
Note that this result generalizes the contraction-reduction lemma (see Lemma A.1) (Feldman et al. 2018) and its variants (Altschuler and Talwar 2022; Altschuler et al. 2024) in existing literature, which all rely on the convexity of the loss function to ensure that the update function is contractive. [^3] A function is said to be contractive if it is $1$-Lipschitz. It characterizes the privacy dynamics of shifted Rényi divergence for noisy stochastic updates with general smooth loss functions. Based on this building block, we are ready to present the privacy guarantees for DPSGD-GC and DPSGD-DC, respectively.
@@
Comparing Theorem 3.4 to 3.3, we observe that the privacy loss for DGSGD-DC can still converge to a constant for non-convex smooth loss function when the model-parameter domain is bounded by $D$. However, unlike the convex cases (Altschuler and Talwar 2022; Altschuler et al. 2024), where the upper bound scales linearly with $D$, the non-convex setting yields a bound that scales quadratically with $D$. This suggests that upper bounds on RDP for bounded-domain scenarios are inherently looser without the assumption on convexity, which is consistent with intuition and recent findings (Kong and Ribero 2024; Chien and Li 2024).
@@ 3.2 Utility Analysis of DPSGD
(Utility bound for DPSGD-GC (Koloskova et al. 2023)). Assume that the population risk function $l(\cdot)$ has smoothness parameter $L$ and SGD variance bounded from above by $\sigma_{\textup{SGD}}^{2}$ (Assumption 3.6). The DPSGD-GC running for $T$ iterations with step-size $\eta\leq\frac{1}{9L}$ has the minimum expected norm of the population risk’s gradient upper bounded by
@@
Unlike previous works on DPSGD-DC (Altschuler and Talwar 2022; Altschuler et al. 2024; Kong and Ribero 2024; Chien and Li 2024), which mainly focused on RDP analysis, our results explicitly demonstrate how gradient clipping and projection affect the utility of DPSGD-DC. As seen from Theorem 3.4, we establish an eventually constant upper bound on privacy loss dependent on the bounded domain diameter $D$, which yields tighter privacy guarantees with smaller $D$. Meanwhile, Proposition 3.12 demonstrates that when the conditions in Theorem 3.10 are satisfied, a smaller $D$ also leads to a lower upper bound on the utility, thus enhancing the privacy-utility trade-off. [^4] A brief discussion on selecting $D$ is provided in Appendix E.3.
@@ 4 Experiment Evaluation
In this section, we present empirical results estimating the privacy level via MIA. Following Kairouz et al. (2015), we estimate $(\epsilon,\delta)$-DP using the false positive rate (FPR) and false negative rate (FNR) of its attack model on the test data, applying the following formula:
@@ A.1 Privacy Amplification by Iteration
Lemma A.1 (Contraction-reduction (Feldman et al. 2018)).
@@
Altschuler et al. (Altschuler and Talwar 2022; Altschuler et al. 2024) provide a mild generalization of this lemma, extending it to be suitable for random contraction maps.
@@ A.3 Lemmas for Privacy Analysis
Definition A.7 ($\infty$-Wasserstein distance (Villani and others 2009)).
@@
Lemma A.8 (Equivalent definitions of $\infty$-Wasserstein distance (Feldman et al. 2018)).
@@
Lemma A.9 (Post-processing inequality (Van Erven and Harremos 2014)).
@@
Lemma A.10 (Partial convexity in the second argument (Van Erven and Harremos 2014)).
@@
Lemma A.11 (Strong composition for Rényi divergence (Mironov 2017)).
@@
Lemma A.13 (Upper bound on Rényi divergence of the SGM (Mironov et al. 2019)).